using the tables at the back of the book, add limit dimensions for the machined hole and shaft, for RC8 fits nominal diameter of 78mm using shaft basis tolerancing.41 15.75 +6.00 6.00 3.50 VALUES IN MILLIMETERS Class RC2 Sliding Fit Nominal Size Range Millimeters Class RC1 Precision Slidi
Answer #1

The data provided in the given question is as follows:

RC8 fit with nominal diameter 78mm

Now the given nominal diameter (78mm) lies in the nominal size range of 50-80

In this range for RC8(loose running) fit the data we have is:

Hole tolerance (H10) = 0.110mm

Shaft tolerance (e9) = 0.080mm

Minimum clearance = 0.150mm

Hole is H10 Shaft is e9. For this notation hole will be represented on the zeroline and shaft will be below the zero line.

Minimum clearance is known which is the difference between hole lower limit and shaft upper limit. So, this value gives the fundamental deviation of the shaft below the zero line.

From all this data we can determine the limits of the hole and shaft as follows:

Hole lower limit = 78.000mm (basic size/zero line)

Hole upper limit = 78 + 0.110 = 78.110mm

Shaft upper limit = 78 - 0.150 = 77.850mm

Shaft lower limit = 77.85 - 0.080 = 77.770mm

Н\o Lollomm ots Zero line 78m) Min-clearance (0.150mm) eq 0.080 mm